A ball dropped from a point P crosses a point Q in t sec. The time taken by it to travel from Q to R,if PQ=QR

Grade:11

1 Answers

Saswat Pani
6 years ago
As the ball is dropped from a point the initial velocity is 0Let the dist PQ=QR=SFrom P to QS=0.5*g*t^2t=√(2S÷g)From P to R2S=0.5*g*t2^2t2=√(4S÷g)t2=√2tSoTime taken from Q to R√2t-t = t(√2-1)